# Binding and spontaneous condensation of excitons in narrow-gap carbon nanotubes

### v1

@@ Acknowledgements
We are grateful to Vikram Deshpande, Assaf Hamo, Shahal Ilani, Ethan Minot, Catalin Spataru, and Gary Steel for inspiring discussions along the years. This work was partly supported by: the MaX – MAterials design at the eXascale – European Centre of Excellence, co-funded by the European High Performance Computing joint Undertaking (JU) and participating countries within the program HORIZON-EUROHPC-JU-2021-COE-1 (Grant No. 101093374); ICSC – Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing – funded by the European Union through the Italian Ministry of University and Research under PNRR M4C2I1.4 (Grant No. CN00000013). We acknowledge ISCRA and ICSC for awarding this project access to the LEONARDO supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CINECA (Italy).
Appendix A End Matter
This End Matter provides additional information on the methodology used in the main text, as it follows. Section A.1 recalls the main results for the accurate screening model derived in Ref. 72 and referred to in Fig. 1(a) through the ‘CNT’ label. Section A.2 details the self-energy correction $\Sigma(k)$ implemented in the BSE calculation. Section A.3 explains the self-consistent calculation of interband coherence, $\Delta(k)$, which is the order parameter of the excitonic insulator phase.
@@ A.1 Screening model
In this section, we summarize the CNT screening model presented in Ref.  72, used to compute the screened electron-hole interaction, $W$. The main novelty is the inclusion of the local field corrections into the Coulomb potential, which originate from the full three-dimensional topology of the Bloch states, $\Psi_{\alpha\tau k}(\mathbf{r})$. In contrast, in the standard effective mass (EM) treatment the Bloch states are graphene-like, i.e, two dimensional. This we accomplish by working with periodically repeated cylindrical supercells. We fix the total length of the tube along the axial direction as $A=N\,a\,\cos\left(\pi/6-\theta\right)$, with $a=2.46$ being the graphene lattice constant. In the non-periodic directions, the supercells are arranged as a square super-lattice of side $\mathfrak{R}$. In order to prevent any superposition between wave functions of different replicas, we set $\mathfrak{R}=7R$, where $R$ is the radius of the nanotube. The reciprocal lattice vectors of the resulting three-dimensional lattice are given by:
$\bm{G}_{\perp}=\frac{\pi}{\mathfrak{R}}\left(n_{1}\hat{x}+n_{3}\hat{z}\right),\quad\bm{G}_{\parallel}=\frac{2\pi}{a\cos\left(\pi/6-\theta\right)}n_{2}\hat{y}.$ (5)
The Bloch states $\Psi$ are expanded over the $\mathbf{G}$ vectors of the three-dimensional supercell lattice. This allows to compute the Coulomb potential $W(\bm{r},\bm{r^{\prime}})$ of the tube by expanding it over a plane-wave basis set:
$W(\bm{r},\bm{r^{\prime}}\!)$ $=$ $\sum_{\bm{q},\bm{G},\bm{G^{\prime}}}\!\!\!\!e^{-i(\bm{G^{\prime}}+\bm{q})\cdot\bm{r^{\prime}}}\!e^{i(\bm{G}+\bm{q})\cdot\bm{r}}\,\epsilon^{-1}_{\bm{G},\bm{G^{\prime}}}(\bm{q})$
×$\quad\times\quad v(\bm{q}+\bm{G^{\prime}}),$
where $\epsilon^{-1}_{\bm{G},\bm{G^{\prime}}}(\bm{q})$ is the inverse dielectric function and $v(\bm{q}+\bm{G^{\prime}})$ is the Fourier transform of the bare Coulomb potential. We prevent any spurious Coulomb interaction between replicas by truncating the bare Coulomb potential along the non-periodic directions at the supercells borders, as detailed in Ref. [91]. The Fourier transform of the truncated Coulomb potential is:
@@
where $J_{0}(x),J_{1}(x)$ are the Bessel functions of first kind, and $K_{0}(x),K_{1}(x)$ the modified Bessel function of second kind. As the Fourier component (7) decreases with the magnitude of the reciprocal lattice vector, the Coulomb potential (A.1) converges with a limited number of $\mathbf{G}$ vectors. We include just the first $G_{\parallel}$ vector $|n_{2}|\leq 1$ and a number of $G_{\perp}$ vectors between $-15\leq n_{1},n_{3}\leq 15$.
@@
$\epsilon_{\bm{G},\bm{G^{\prime}}}(\bm{q})=\delta_{\bm{G},\bm{G^{\prime}}}-\Pi_{\bm{G},\bm{G^{\prime}}}(q)\,v(\bm{q}+\bm{G}),$ (8)
@@
$\Pi_{\bm{G},\bm{G^{\prime}}}(q)\quad=\quad\Pi^{\text{NT}}_{\bm{G},\bm{G^{\prime}}}(q)\\ \times\quad\left\{5\cos\!\left[2.7\left(\pi/6-\theta\right)\!\right]\!Rq+3.806\,[R/(\text{1 nm})]^{1.46}\right\}.$ (9)
@@
$\Pi^{\text{NT}}_{\bm{G},\bm{G^{\prime}}}(q)=-\frac{2}{\pi\gamma}J_{0}(RG_{\perp})J_{0}(RG^{\prime}_{\perp})\\ \sum_{\tau}\Bigg{[}1+\frac{2k_{\tau}^{2}}{q\sqrt{q^{2}+4k_{\tau}^{2}}}\log\Bigg(\frac{\sqrt{q^{2}+4k_{\tau}^{2}}-q}{\sqrt{q^{2}+4k_{\tau}^{2}}+q}\Bigg{missing})\Bigg{]}$ (10)
is modified according to (9), with the inclusion of a $q-$ and $\theta$-dependent multiplicative correction factor that slighlty differs from unity. This multiplicative factor has been determined fitting first-principles results, specifically for the $G$ components of $\epsilon^{-1}_{\bm{G},\bm{G^{\prime}}}(q)$ (details can be found in Ref. 72).
The interaction matrix element associated with electron-hole binding is calculated by projecting the screened potential, $W(\bm{r},\bm{r^{\prime}})$, from an initial $e$-$h$ state $(c,\tau,k)(v,\tau,k+q)$ to a final state $(v,\tau,k)(c,\tau,k+q)$ within the same valley $\tau$:
$W^{\tau}(k,k+q)=\sum_{\bm{G}}\sum_{\bm{G^{\prime}}}\langle c\tau k|e^{-i(\bm{G^{\prime}}+\bm{q})\cdot\bm{r^{\prime}}}|c\tau k+q\rangle$
$\langle v\tau k+q|e^{i(\bm{G}+\bm{q})\cdot\bm{r}}|v\tau k\rangle\ \epsilon^{-1}_{\bm{G},\bm{G^{\prime}}}(q)v(\bm{q}+\bm{G^{\prime}}).$ (11)
@@
$W^{\tau}(k,k+q)=(\bm{F}^{\tau\dagger}_{vk}\bm{F}^{\tau}_{vk+q})(\bm{F}^{\tau\dagger}_{ck+q}\bm{F}^{\tau}_{ck})\,W(q),$ (12)
@@
$\!\!W(q)\!=\sum_{\bm{G},\bm{G^{\prime}}}\!\!J_{0}(RG_{\perp})J_{0}(RG^{\prime}_{\perp})\,\epsilon^{-1}_{\bm{G},\bm{G^{\prime}}}(q)\,v(\bm{q}+\bm{G^{\prime}}).$ (13)
@@ A.2 Self-energy correction
[…]
@@
[…]
@@
$G_{0}(\bm{r},\bm{r}^{\prime},\omega)=\!\!\sum_{\alpha,\tau,k}\!\frac{\Psi^{*}_{\alpha\tau k}(\bm{r})\Psi_{\alpha\tau k}(\bm{r^{\prime}})}{\omega-\varepsilon_{\alpha}(k)\!+i\eta\ \textrm{sign}[\varepsilon_{\alpha}(k)]},$ (17)
@@
$\Sigma_{\alpha}(k)=-\frac{1}{3A}\sum_{q}|\bm{F}^{\tau\dagger}_{\alpha k}\bm{F}^{\tau}_{\alpha k+q}|^{2}W(q),$ (18)
@@
$\Sigma(k)=\Sigma_{c}(k)-\Sigma_{v}(k)=\\ \frac{1}{3A}\sum_{q}(|\bm{F}^{\tau\dagger}_{vk}\bm{F}^{\tau}_{vk+q}|^{2}-|\bm{F}^{\tau\dagger}_{ck}\bm{F}^{\tau}_{ck+q}|^{2})\,W(q).$ (19)
@@ A.3 Interband coherence in the excitonic insulator phase
$\mathcal{H}=\sum_{\tau k\sigma}E^{0}_{\tau k}\left(c^{\tau\dagger}_{k\sigma}c^{\tau}_{k\sigma}-v^{\tau\dagger}_{k\sigma}v^{\tau}_{k\sigma}\right)\\ -\frac{1}{A}\sum_{\tau kk^{\prime}}\sum_{\sigma\sigma^{\prime}}W^{\tau}(k,k^{\prime})\ c^{\tau\dagger}_{k\sigma}v^{\tau}_{k\sigma^{\prime}}v^{\tau\dagger}_{k^{\prime}\sigma^{\prime}}c^{\tau}_{k^{\prime}\sigma}\\ -\frac{1}{A}\sum_{\tau\neq\tau^{\prime}}\sum_{kk^{\prime}}\sum_{\sigma\sigma^{\prime}}W^{\tau\tau^{\prime}}c^{\tau\dagger}_{k\sigma^{\prime}}v^{\tau}_{k\sigma}v^{\tau^{\prime}\dagger}_{k^{\prime}\sigma^{\prime}}c^{\tau^{\prime}}_{k^{\prime}\sigma}.$ (21)
Here $E^{0}_{\tau k}=\gamma\left[k_{c}^{2}+k^{2}\right]^{1/2}+\Sigma(k)/2$ is the energy of conduction-band quasiparticles. Whereas the sum over $k$ and $k^{\prime}$ in principle extends throughout the Brillouin zone, practically one may limit it to $-0.1\times 2\pi/a<k<0.1\times 2\pi/a$, due to the large Fermi velocity of the Dirac cone.
The energy of the noninteracting band insulator, $\mathcal{E}_{0}$, is:
@@
$\mathcal{E}_{\text{EI}}-\mathcal{E}_{0}=2\sum_{\tau k}E^{0}_{\tau k}(v_{\tau k}^{2}-u_{\tau k}^{2}+1)\\ -\frac{2}{A}\sum_{\tau kk^{\prime}}W^{\tau}(k,k^{\prime})v_{\tau k}u_{\tau k}v_{\tau k^{\prime}}u_{\tau k^{\prime}}\\ -\frac{2}{A}\sum_{\tau\neq\tau^{\prime}}\sum_{kk^{\prime}}W^{\tau\tau^{\prime}}v_{\tau k}u_{\tau k}v_{\tau^{\prime}k^{\prime}}u_{\tau^{\prime}k^{\prime}}.$ (23)
@@
$\mathcal{E}_{\text{EI}}-\mathcal{E}_{0}=2\sum_{\tau k}\left[E^{0}_{\tau k}(v_{\tau k}^{2}-u_{\tau k}^{2}+1)-\left|\Delta(\tau k)\right|u_{\tau k}v_{\tau k}\right].$ (25)
@@
$u_{\tau k}^{2}=$ 12($\ \frac{1}{2}\left(1+\frac{E^{0}_{\tau k}}{2E_{\tau k}}\right),$ (26)

### v2

@@ Acknowledgements
We are grateful to Vikram Deshpande, Assaf Hamo, Shahal Ilani, Ethan Minot, Catalin Spataru, and Gary Steel for inspiring discussions along the years. This work was partly supported by: the MaX – MAterials design at the eXascale – European Centre of Excellence, co-funded by the European High Performance Computing joint Undertaking (JU) and participating countries within the program HORIZON-EUROHPC-JU-2021-COE-1 (Grant No. 101093374); ICSC – Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing – funded by the European Union through the Italian Ministry of University and Research under PNRR M4C2I1.4 (Grant No. CN00000013).
We acknowledge ISCRA and ICSC for awarding this project access to the LEONARDO supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CINECA (Italy).
Appendix A
This Appendix provides additional information on the methodology used in the main text, as it follows. Section A.1 recalls the main results for the accurate screening model derived in Ref. 73 and referred to in Fig. 1(a) through the ‘CNT’ label. Section A.2 details the self-energy correction $\Sigma(k)$ implemented in the BSE calculation. Section A.3 explains the self-consistent calculation of interband coherence, $\Delta(k)$, which is the order parameter of the excitonic insulator phase.
@@ A.1 Screening model
In this section, we summarize the CNT screening model presented in Ref.  73, used to compute the screened electron-hole interaction, $W$. The main novelty is the inclusion of the local field corrections into the Coulomb potential, which originate from the full three-dimensional topology of the Bloch states, $\Psi_{\alpha\tau k}(\mathbf{r})$. In contrast, in the standard effective mass (EM) treatment the Bloch states are graphene-like, i.e, two dimensional. This we accomplish by working with periodically repeated cylindrical supercells. We fix the total length of the tube along the axial direction as $A=N\,a\,\cos\left(\pi/6-\theta\right)$, with $a=2.46$ Å being the graphene lattice constant. In the non-periodic directions, the supercells are arranged as a square super-lattice of side $\mathfrak{R}$. In order to prevent any superposition between wave functions of different replicas, we set $\mathfrak{R}=7R$, where $R$ is the radius of the nanotube. The reciprocal lattice vectors of the resulting three-dimensional lattice are given by:
$\boldsymbol{G}_{\perp}=\frac{\pi}{\mathfrak{R}}\left(n_{1}\hat{x}+n_{3}\hat{z}\right),\quad\boldsymbol{G}_{\parallel}=\frac{2\pi}{a\cos\left(\pi/6-\theta\right)}n_{2}\hat{y}$. (5)
The Bloch states $\Psi$ are expanded over the $\mathbf{G}$ vectors of the three-dimensional supercell lattice. This allows to compute the Coulomb potential $W(\boldsymbol{r},\boldsymbol{r^{\prime}})$ of the tube by expanding it over a plane-wave basis set:
$W(\boldsymbol{r},\boldsymbol{r^{\prime}}\!)$ $=$ $\sum_{\boldsymbol{q},\boldsymbol{G},\boldsymbol{G^{\prime}}}\!\!\!\!e^{-i(\boldsymbol{G^{\prime}}+\boldsymbol{q})\cdot\boldsymbol{r^{\prime}}}\!e^{i(\boldsymbol{G}+\boldsymbol{q})\cdot\boldsymbol{r}}\,\epsilon^{-1}_{\boldsymbol{G},\boldsymbol{G^{\prime}}}(\boldsymbol{q})$
$\quad\times\quad v(\boldsymbol{q}+\boldsymbol{G^{\prime}}),$
where $\epsilon^{-1}_{\boldsymbol{G},\boldsymbol{G^{\prime}}}(\boldsymbol{q})$ is the inverse dielectric function and $v(\boldsymbol{q}+\boldsymbol{G^{\prime}})$ is the Fourier transform of the bare Coulomb potential. We prevent any spurious Coulomb interaction between replicas by truncating the bare Coulomb potential along the non-periodic directions at the supercells borders, as detailed in Ref. [92]. The Fourier transform of the truncated Coulomb potential is:
@@
where $J_{0}(x),J_{1}(x)$ are the Bessel functions of first kind, and $K_{0}(x),K_{1}(x)$ the modified Bessel function of second kind. As the Fourier component (7) decreases with the magnitude of the reciprocal lattice vector, the Coulomb potential (A.1) converges with a limited number of $\mathbf{G}$ vectors. We include just the first $G_{\parallel}$ vector $|n_{2}|\leq 1$ and a number of $G_{\perp}$ vectors between $15\leq n_{1},n_{3}\leq 15$.
@@
$\epsilon_{\boldsymbol{G},\boldsymbol{G^{\prime}}}(\boldsymbol{q})=\delta_{\boldsymbol{G},\boldsymbol{G^{\prime}}}-\Pi_{\boldsymbol{G},\boldsymbol{G^{\prime}}}(q)\,v(\boldsymbol{q}+\boldsymbol{G}),$ (8)
@@
$\Pi_{\boldsymbol{G},\boldsymbol{G^{\prime}}}(q)\quad=\quad\Pi^{\text{NT}}_{\boldsymbol{G},\boldsymbol{G^{\prime}}}(q)\\ \times\quad\left\{5\cos\!\left[2.7\left(\pi/6-\theta\right)\!\right]\!Rq+3.806\,[R/(\text{1 nm})]^{1.46}\right\}.$ (9)
@@
[…]
is modified according to (9), with the inclusion of a $q$- and $\theta$-dependent multiplicative correction factor that slighlty differs from unity. This multiplicative factor has been determined fitting first-principles results, specifically for the $G$ components of $\epsilon^{-1}_{\boldsymbol{G},\boldsymbol{G^{\prime}}}(q)$ (details can be found in Ref. 73).
The interaction matrix element associated with electron-hole binding is calculated by projecting the screened potential, $W(\boldsymbol{r},\boldsymbol{r^{\prime}})$, from an initial $e$-$h$ state $(c,\tau,k)(v,\tau,k+q)$ to a final state $(v,\tau,k)(c,\tau,k+q)$ within the same valley $\tau$:
$W^{\tau}(k,k+q)=\sum_{\boldsymbol{G}}\sum_{\boldsymbol{G^{\prime}}}\langle c\tau k|e^{-i(\boldsymbol{G^{\prime}}+\boldsymbol{q})\cdot\boldsymbol{r^{\prime}}}|c\tau k+q\rangle$
$\langle v\tau k+q|e^{i(\boldsymbol{G}+\boldsymbol{q})\cdot\boldsymbol{r}}|v\tau k\rangle\ \epsilon^{-1}_{\boldsymbol{G},\boldsymbol{G^{\prime}}}(q)v(\boldsymbol{q}+\boldsymbol{G^{\prime}}).$ (11)
@@
$W^{\tau}(k,k+q)=(\boldsymbol{F}^{\tau\dagger}_{vk}\boldsymbol{F}^{\tau}_{vk+q})(\boldsymbol{F}^{\tau\dagger}_{ck+q}\boldsymbol{F}^{\tau}_{ck})\,W(q),$ (12)
@@
$\!\!W(q)\!=\sum_{\boldsymbol{G},\boldsymbol{G^{\prime}}}\!\!J_{0}(RG_{\perp})J_{0}(RG^{\prime}_{\perp})\,\epsilon^{-1}_{\boldsymbol{G},\boldsymbol{G^{\prime}}}(q)\,v(\boldsymbol{q}+\boldsymbol{G^{\prime}}).$ (13)
@@ A.2 Self-energy correction
$\Sigma_{\alpha\tau}(k)=B\!\int\!\!\!\int\!\!d\mathbf{r}\,d\mathbf{r^{\prime}}\Psi_{\alpha\tau k}(\boldsymbol{r})\,\Sigma^{HF}(\boldsymbol{r},\boldsymbol{r}^{\prime})\,\Psi^{*}_{\alpha\tau k}(\boldsymbol{r^{\prime}}),$ (15)
@@
$\Sigma^{HF}(\boldsymbol{r},\boldsymbol{r}^{\prime})=\lim_{\eta\rightarrow 0^{+}}\frac{i}{2\pi}\int\!\!d\omega^{\prime}G_{0}(\boldsymbol{r},\boldsymbol{r}^{\prime},\omega^{\prime})W(\boldsymbol{r},\boldsymbol{r}^{\prime})e^{i\eta\omega^{\prime}},$ (16)
@@
$G_{0}(\boldsymbol{r},\boldsymbol{r}^{\prime},\omega)=\!\!\sum_{\alpha,\tau,k}\!\frac{\Psi^{*}_{\alpha\tau k}(\boldsymbol{r})\Psi_{\alpha\tau k}(\boldsymbol{r^{\prime}})}{\omega-\varepsilon_{\alpha}(k)\!+i\eta\ \textrm{sign}[\varepsilon_{\alpha}(k)]},$ (17)
@@
$\Sigma_{\alpha}(k)=-\frac{1}{3A}\sum_{q}|\boldsymbol{F}^{\tau\dagger}_{\alpha k}\boldsymbol{F}^{\tau}_{\alpha k+q}|^{2}W(q),$ (18)
@@
$\Sigma(k)=\Sigma_{c}(k)-\Sigma_{v}(k)=\\ \frac{1}{3A}\sum_{q}(|\boldsymbol{F}^{\tau\dagger}_{vk}\boldsymbol{F}^{\tau}_{vk+q}|^{2}-|\boldsymbol{F}^{\tau\dagger}_{ck}\boldsymbol{F}^{\tau}_{ck+q}|^{2})\,W(q).$ (19)
@@ A.3 Interband coherence in the excitonic insulator phase
$\mathcal{H}=\sum_{\tau k\sigma}E^{0}_{\tau k}\left(c^{\tau\dagger}_{k\sigma}c^{\tau}_{k\sigma}-v^{\tau\dagger}_{k\sigma}v^{\tau}_{k\sigma}\right)\\ -\frac{1}{A}\sum_{\tau kk^{\prime}}\sum_{\sigma\sigma^{\prime}}W^{\tau}(k,k^{\prime})\ \left(c^{\tau\dagger}_{k\sigma}v^{\tau}_{k\sigma^{\prime}}v^{\tau\dagger}_{k^{\prime}\sigma^{\prime}}c^{\tau}_{k^{\prime}\sigma}+c.c.\right)\\ -\frac{1}{A}\sum_{\tau\neq\tau^{\prime}}\sum_{kk^{\prime}}\sum_{\sigma\sigma^{\prime}}W^{\tau\tau^{\prime}}\left(c^{\tau\dagger}_{k\sigma^{\prime}}v^{\tau}_{k\sigma}v^{\tau^{\prime}\dagger}_{k^{\prime}\sigma^{\prime}}c^{\tau^{\prime}}_{k^{\prime}\sigma}+c.c.\right).$ (21)
Here $E^{0}_{\tau k}=\gamma\left[k_{c}^{2}+k^{2}\right]^{1/2}+\Sigma(k)/2$ is the energy of conduction-band quasiparticles. Whereas the sum over $k$ and $k^{\prime}$ in principle extends throughout the Brillouin zone, practically one may limit it to $0.1\times 2\pi/a<k<0.1\times 2\pi/a$, due to the large Fermi velocity of the Dirac cone.
The energy of the non-interacting band insulator, $\mathcal{E}_{0}$, is:
@@
$\mathcal{E}_{\text{EI}}-\mathcal{E}_{0}=2\sum_{\tau k}E^{0}_{\tau k}(v_{\tau k}^{2}-u_{\tau k}^{2}+1)\\ -\frac{4}{A}\sum_{\tau kk^{\prime}}W^{\tau}(k,k^{\prime})v_{\tau k}u_{\tau k}v_{\tau k^{\prime}}u_{\tau k^{\prime}}\\ -\frac{4}{A}\sum_{\tau\neq\tau^{\prime}}\sum_{kk^{\prime}}W^{\tau\tau^{\prime}}v_{\tau k}u_{\tau k}v_{\tau^{\prime}k^{\prime}}u_{\tau^{\prime}k^{\prime}}.$ (23)
@@
$\mathcal{E}_{\text{EI}}-\mathcal{E}_{0}=2\sum_{\tau k}\left[E^{0}_{\tau k}(v_{\tau k}^{2}-u_{\tau k}^{2}+1)-2\left|\Delta(\tau k)\right|u_{\tau k}v_{\tau k}\right].$ (25)
@@
$u_{\tau k}^{2}=$ $\ \frac{1}{2}\left(1+\frac{E^{0}_{\tau k}}{2E_{\tau k}}\right),$ (26)